OpenStudy (anonymous):

Maria is working at a local retail store. She is making $5 per hour plus a 5% commissionon the sales she makes. How much money would Maria make if she worked 10 hours and sold $2,000 worth of merchandise? A. $100 B. $150 C. $50 D. 250

OpenStudy (nubeer):

Earn = 5H +(5/100)*2000 H is the number of hours.

OpenStudy (anonymous):

h=60

OpenStudy (nubeer):

no.. look in your question it says she work for 10 hours.. so H = 10

OpenStudy (anonymous):

oh okay hold on

OpenStudy (anonymous):

5x10+60=---------> 50+60=110

OpenStudy (anonymous):

wow thats not even an answer up there .. lol

OpenStudy (nubeer):

no.. u are wrong at the 60 part..

OpenStudy (nubeer):

5/100 =?

OpenStudy (anonymous):

20

OpenStudy (nubeer):

nooo..\[\frac{ 5 }{ 100 } =?\]

OpenStudy (anonymous):

0.05

OpenStudy (nubeer):

yes good.. now 0.05 * 2000 = ?

OpenStudy (anonymous):

100

OpenStudy (nubeer):

good :) now 5*10 + 100 =?

OpenStudy (anonymous):

150

OpenStudy (nubeer):

there u go.. u got it :)

OpenStudy (anonymous):

thank you :) life saver for sure !:)

OpenStudy (nubeer):

hahaha .. you're welcome :)